HUBBLE SPACE TELESCOPE DAILY REPORT #2134 PERIOD COVERED: 0000Z (UTC) 05/29/98 - 0000Z (UTC) 06/01/98 Daily Status Report as of 152/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7282 (A Continuing Snapshot Survey for Compact Stellar Groups and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make observations of the Wolf-Rayet star WR80 as part of a continuing snapshot survey of compact stellar groups and clusters. The newly imaged clusters allow us to accurately determine the distances, luminosities, and ages of the WR members. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-3865 and NGC-5064 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.5 Completed Two Sets of WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star cluster NGC-1831 which is located within the Large Magellanic Cloud in order to study its formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.6 Completed WF/PC-2 7470 (A Snapshot Survey of Galactic Globular Clusters) The WF/PC-2 was used to make snap shot survey observations of the crowded center of the galactic globular cluster NGC-6139. The resulting color-magnitude diagram will include all stars down to about one magnitude below the main-sequence turnoff. The observations were completed as scheduled, and no problems were reported. 1.7 Completed Three Sets of NIC/1/2/3 7372 (Fundamental Plane, Morphology-Density Relation, and Lensing in the z=0.83 Cluster) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galactic cluster Gal-Clus-1. These data will be used to augment the HST mosaics of the clusters with extensive ground-based spectroscopy and imaging at the Keck, MMT, NTT, and WHT telescopes. The observations were completed as planned, and no problems were reported. 1.8 Completed Ten Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. During the eighth iteration of this proposal, the loss of lock described in 2.1 occurred, temporarily taking down the take data flag. Otherwise, the observations were completed as planned, and no further problems were reported. 1.9 Completed WF/PC-2 7590 (Cosmological Parameters Omega and Lambda From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova 11H0.9Z2 in order to study its redshift. The HST was used to make direct measurements of the cosmological parameters Omega, Lambda, and curvature using Type 1A supernovae as a calibrated "standard candle". The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.11 Completed NIC/1/2/3 7902 (Photometric Monitoring, Photometric Monitoring-II) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of absolute standard, solar analog star P330E-Cal as part of a photometric monitoring program. Observations were made at several different wavelengths spanning the NICMOS wavelength range in each camera NIC/1, NIC/2, and NIC/3 to monitor the photometric stability of the NICMOS cameras as a function of wavelength. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/1/2 7861 (A Search for Extrasolar Giant Planets in the Nearby TW Hya Association) The Near Infrared Camera (Cameras One and Two) was used to search for extrasolar giant planets on the nearby star CD-298887 (a.k.a., TW Hya). One of the primary goals of NASA's mission is the identification of formed and nascent planetary systems. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.14 Completed NIC/3 7830 (IR Spectra for Known-Mass M Dwarfs) The Near Infrared Camera (Camera Three) was used to make infrared spectrographic observations of M-Class dwarf star GJ473AB. Over the last decade there has been considerable debate on two major issues in the low-mass star literature: {1} the effective temperature scale for M dwarfs and its relation to mass and luminosity and {2} the conversion from an absolute magnitude in some passband to a bolometric magnitude. As detailed in 2.1, the acquisition for this proposal defaulted to fine lock backup on one FGS, possibly affecting the five abservations contained therein. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.15 Completed NIC/2 7887 (A Survey of Gravitational Lenses as Cosmological Tools II) The Near Infrared Camera (Camera Two) was used to make observations of the gravitational lens candidate system Q0957+561. The propose is to extend our IR survey of gravitational lenses and gravitational lens candidates by obtaining NICMOS images of 14 additional lens systems or candidates. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Seven Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.17 Completed WF/PC-2 7277 (Calibration of the Cepheid P-L Relation with Observations of the Maser-Host Galaxy NGC) The WF/PC-2 was used to make observations of the maser-host galaxy NGC-4258 to study its Cepheid variable stars. These data will help refine the calibration of the Cepheid period-luminosity relationship. The observations were completed as scheduled, and no problems were reported. 1.18 Completed WF/PC-2 7409 (Resolving sdB Binary Systems) The WF/PC-2 was used to make observations to try and resolve the sdB class binary system STAR-09451. This is a continuation of the approved Cycle 6 Proposal No. 6351, which due to unlucky circumstances did not succeed in obtaining any data until 8/30/96. The observations were completed as scheduled, and no anomalies were reported. 1.19 Completed NIC/1 7873 (NICMOS Observations of Potential JVAS/CLASS Gravitational Lenses) The Near Infrared Camera (Camera One) was used to make observations of the potential gravitational lens system J2321+0527. The JVAS/CLASS surveys have been the most successful to date at finding arcsecond-scale gravitational lenses, almost doubling the known number {11 lenses found of which 10 are new discoveries}. The observations were executed as scheduled, and no anomalies were noted. 1.20 Completed NIC/1/2/3 7883 (Infrared Imaging of High Redshift {0.4 <= z <= 1} Tully-Fisher Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the high redshifted Tully- Fisher class galaxy SA68-KK687. With the advent of 10 meter class telescopes combined with the high spatial resolution of the HST, it is now possible to study the internal kinematics and structural properties of galaxies out to a redshift of z ~ 1. During this proposal, the loss of lock described in 2.1 occurred, temporarily taking down the take data flag. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.21 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.22 Completed STIS/CCD 7672 (STIS CCD Sensitivity Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star AGK+81D266 in order to monitor the sensitivity of each CCD grating mode to detect any change due to contamination or other causes. The observations were completed as scheduled, and no anomalies were noted. 1.23 Completed NIC/1/2 7864 (The Host Galaxies of Radio-Quiet High- Redshift Quasars) The Near Infrared Camera (Cameras One and Two) was used to make observations of the radio-quite, high red shifted quasar MZZ9592 and its host galaxy. The underlying causes of the rapid cosmological evolution of the quasar population are a mystery around which are woven many different strands of contemporary cosmology. The observations were executed as scheduled, and no anomalies were noted. 1.24 Completed FGS 7493 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) The FGSs were used to make observations of star GJ2005 in order to make calibration observations of the mass-luninocity relation at the end of the main sequence. The study utilizes HST-FGS and infrared speckle imaging to calibrate the mass-luminosity relation for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. The observations were completed as scheduled, and no problems were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 22 Successful : 22 The acquisition at 150/050218Z defaulted to fine lock backup on FGS-2 only when FGS-1 exceeded the scan step limit. HSTAR 6497 was written. The proposal described in 1.14 may have been affected. As described in HSTAR 6498, there was a loss of lock at 151/000856Z. The proposals mentioned in 1.8 and 1.20 were affacted. Scheduled Reacquisitions : 21 Successful : 21 2.2 FHST Updates: Scheduled : 44 Successful : 44 2.3 Operations Notes: The SSR EDAC error counter was cleared seven times using ROP NS-1. As documented in HSTAR 6499, the parameter NSBITERR (the NICMOS EDAC single bit error count) flagged out-of-limits low at 151/134856Z. Per direction from an operations note, the STScI was notified, the limits was raised by one count, and ROP NS-11 (Part A) was executed to dump and clear the NICMOS error log. This was accomplished at 151/1653Z. The STIS MCE-1 reset at 152/0916Z while in high voltage and while outside an SAA contour. The MCE-1 will be recovered via the SMS at the time of the next MAMA-1 turn-on at 157/054412Z. HSTAR 6500 documents this event. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH